OpenStudy (anonymous):

Given x=(1,1,1,1)^T and y =(8,2,2,0)^T b) Find the vector projection of p of x onto y

OpenStudy (zarkon):

what is 'p of x'

OpenStudy (zarkon):

are you just projecting x onto y?

OpenStudy (anonymous):

yes, projecting x onto y

OpenStudy (zarkon):

\[\frac{x\cdot y}{y\cdot y}y\]

OpenStudy (zarkon):

so do the dot product of x and y then the dot product of y and y divide the two...then multiply by y

OpenStudy (anonymous):

ok, it is just there is just so many formula to memorize in linear algebra

OpenStudy (anonymous):

thanks so much